OpenStudy (anonymous):

f(x) = x/(x-3) please state how to find the domain of this problem thanks so much

OpenStudy (ericwise):

Find the non-permissible values of the denominator, ie the denominator cannot be zero, this can be represented by the following (x-3)=0 *not equal to zero*

OpenStudy (ericwise):

thus, the NPV is +3

OpenStudy (ericwise):

So the domain is: {X e R | x =/ 0}

OpenStudy (anonymous):

so would that be all real numbers greater than 0??

OpenStudy (ericwise):

sorry... {X e R | x =/ 3}

OpenStudy (anonymous):

what does X and R stand for in the formula

OpenStudy (ericwise):

no just above zero, but also below

OpenStudy (anonymous):

so the absolute value of x greater and less than 0

OpenStudy (ericwise):

"x" reffering to the variable, the 'e' means "Element of" and the 'R' means "Real Numbers"

OpenStudy (ericwise):

When dealing with absolute values, all results are positive, could you post your question pls?

OpenStudy (anonymous):

so would it be lxl>0

OpenStudy (ericwise):

true

OpenStudy (ericwise):

greater than or equal to**

OpenStudy (anonymous):

okay yeah thanks so lxl≥0

OpenStudy (anonymous):

\[lxl \ge0\]

OpenStudy (ericwise):

then with it, add that x is not equal to 3 and that's your domain. Now notations vary and I may be formatting differently than what your teacher expects

OpenStudy (ericwise):

but that's the gist of it: {|x| >_ 0} U {x =/ 3} That's set builder notation, try interpreting it and convert it into whatever format you use.

OpenStudy (anonymous):

so then wouldnt it be lxl>3, lxl<3 D: (-∞, 3) U (3,∞)

OpenStudy (anonymous):

\[D: (-\infty,3) \cup(3,\infty)\]

OpenStudy (ericwise):

not quite, see, at x=3, there is a vertical seperation between two 'broken' parts of the graph (called a vertical asymptote) this means that x exists at all values (including 2.9999999...) but not at 3

OpenStudy (ericwise):

actually yes, that last re is correct

OpenStudy (anonymous):

isnt that what ")" and not "]" is for

OpenStudy (ericwise):

I was replying to a more recent one, you got it

OpenStudy (anonymous):

thanks

OpenStudy (ericwise):

glad to help ^_^

OpenStudy (ericwise):

have a good one m8

OpenStudy (anonymous):

you too :)